If Gina leaves now and drives at 66 km/h, she will reach Alton just in time for her appointment. On the other hand, if she has lunch first and leaves in 40 minutes, she will have to drive at 90 km/h to make her appointment. How far away is Alton?
Answer by mananth(16946)

let Alton be x miles from where she is.
Time taken in both options is the same.
Option I speed = 60 mph
t=d/r
time taken = x/60
she will take 40 minutes less at higher speed. = 40/60 = 2/3 hours

t-2/3 = x/90
t= x/90 +2/3
Equate the time in both options
x/60 = x/90 +2/3
x/60 -x/90 = 2/3
LCD = 180
Multiply equation by 180
3x-2x=180*2/3
x=120
Alton is 120 miles from where she is now.